OpenStudy (anonymous):

simplify 15x^2/5sqrt(x)

OpenStudy (anonymous):

multiply by sqrt{x}/sqrt{x} 15x^2sqrt{x}/5x 3xsqrt{x}

OpenStudy (anonymous):

3x^3/2

OpenStudy (anonymous):

\[\frac{15(\sqrt{x}\sqrt{x}\sqrt{x}\sqrt{x}))}{5\sqrt{x}}\]